# Exploiting the Structure in Tensor Decompositions for Matrix Multiplication

### v1

@@ Abstract.
conference: ; ; $c$s: Computing methodologies Algebraic algorithms
@@ 1. Introduction
We ran a search for such algorithms, first using a flip graph search to find algorithms using a minimal number of multiplications while retaining the special structure needed to apply our new recursive method. Then we optimize the number of additions using DeGroote actions. These steps produce an algorithm that is valid only modulo $2$, thus we apply Hensel lifting to lift the algorithm to work over the integers, and finally we reduce the number of additions further by Mårtensson and Wagner’s common subexpression elimination method (Mårtensson and Wagner, 2025). We find improvements in the exponent for several small base cases including $3\times 3$ matrix multiplication, where we reduce the exponent from Laderman’s $2.854$ (Laderman, 1976) to $2.836$.
@@ 2. Background
Strassen’s algorithm (STRASSEN, 1969) uses a divide and conquer strategy to multiply matrices faster than the standard algorithm. First, it reduces the multiplication of two $n\times n$ matrices to multiplying two $2\times 2$ matrices whose entries are $\frac{n}{2}\times\frac{n}{2}$ matrices using block matrix multiplication. Then it computes this $2\times 2$ block matrix product using only $7$ multiplications of the subblocks by linearly combining them in a clever way. Using this strategy recursively gives an algorithm to multiply $n\times n$ matrices using $\operatorname{O}(n^{\log_{2}(7)})$ operations in $R$. The exponent only depends on the number of multiplications, since the cost of the recursive call dominates the cost of the linear combinations.
@@
where $\mathbf{A}\in A,\mathbf{B}\in B$, $\mathbf{A}_{ij},\mathbf{B}_{ij}$ represent the $(i,j)$-th entry of $\mathbf{A}$ and $\mathbf{B}$ respectively and $\varGamma_{ij}\in C$ is the matrix with all zeros except for a $1$ at the $(i,j)$-th entry. Note the $c_{ij}$ corresponds to $\mathbf{C}_{ji}$ in the product matrix; this is a standard convention to keep the cyclic nature of matrix multiplication.
@@
Using the decomposition Strassen found, we get $\langle 2,2,2\rangle\leq\langle 7\rangle$ and hence $\omega\leq\log_{2}(7)$. Further analysis allows to compute the leading coefficient of this algorithm as well, as we will see next. While the exponent is given by the rank of a decomposition, i.e. the number of multiplications, the leading coefficient depends on the number of additions and scalar multiplications needed in the restriction. We will denote this number by $A(\phi)$ if the restriction is given by $\phi$. The homomorphism for a restriction is in general not unique, but in our case every restriction is witnessed by an explicit homomorphism, so we just write $A$ when the homomorphism is clear from the context.
@@
We will show by induction on $N$ that
@@
In the case of Strassen’s original algorithm, which used a restriction $\langle 2,2,2\rangle\leq_{\phi}\langle 7\rangle$ with $A(\phi)=18$, we find that this theorem gives us a bound on the leading coefficient $L\leq 40$. Usually the leading coefficient bounds are computed in the idealized case where every recursive call is assumed to be done exactly. In this case, the leading coefficient of Strassen’s algorithm is $7$, the general formula being $L\leq\frac{A}{r-n^{2}}+1$ (Karstadt and Schwartz, 2017).
@@
While we have to compute $\langle 1,1,2\rangle,\langle 1,2,2\rangle$ and their cyclic permutations by the standard algorithm, using $2$ and $4$ multiplications respectively, we can compute $\langle 2,2,2\rangle$ using Strassen’s algorithm with $7$ multiplications. In the language of tensor restrictions we use $\langle 1,1,2\rangle\leq\langle 2\rangle$, $\langle 1,2,2\rangle\leq\langle 4\rangle$ and $\langle 2,2,2\rangle\leq\langle 7\rangle$, to conclude that $\langle 216,216,216\rangle\leq\langle 3581065\rangle$. Applying Theorem 3 to this restriction gives an algorithm with exponent $\omega_{0}=2.80751$, which is only slightly smaller than the original exponent $2.80754$. We can do even better by taking higher tensor powers.
@@ 4. Complexity Analysis
We will prove by induction on $N$ that
@@
We can count the arithmetic cost by considering performing this restriction recursively: we perform $A$ arithmetic operations on matrices of size $n^{k-1}\times n^{k-1}$ at the first step, so the arithmetic operation count at this step is bounded by $An^{2(k-1)}$. At the second step we have at most $r$ lots of $n^{k-2}\times n^{k-2}$ matrices that we need to construct, and then use linear combinations to get outputs; this means the arithmetic operations at this recursive step is bounded by $Arn^{2(k-2)}$. Continuing this, we see that at the $j$-th recursive step, we have at most $Ar^{j-1}n^{2(k-j)}$ arithmetic operations. Summing all of these, we get that the arithmetic cost is bounded by the geometric sum
@@ 5. Specific Decompositions
We have already used the Moosbauer and Poole’s tensor decomposition of $\langle 6,6,6\rangle$ in the form $137\odot\langle 1\rangle\oplus 8\odot\langle 2,1,1\rangle$ as illustrating example. According to Theorem 2, this decomposition leads to an exponent $2.80496$, slightly better than the exponent $2.80754$ announced by Moosbauer and Poole and than Strassen’s exponent $2.80735$, though not as good as the exponent $2.7925$ of the decompositions of Novikov et al. (Novikov et al., 2025b) and Dumas, Pernet, and Sedoglavic (Dumas et al., 2025). We have repeated the computation of Moosbauer and Pool in order to generate further decompositions of $\langle 6,6,6\rangle$ in the hope to find some that contain more copies of $\langle 2,1,1\rangle$. The best we found contains $18$ such copies and thus leads to an exponent $2.8016$. The leading coefficient for this decomposition turns out to be $7.22$, compared to $6$ for Strassen’s algorithm.
@@
In a next step, for various choices of $n,m,p$, we have searched for decompositions of $\langle n,m,p\rangle$ that contain copies of $\langle 1,2,2\rangle$ and/or $\langle k,1,1\rangle$ for $k=2,3,4$, or permutations of these. This search procedure is described in more detail below. The results are summarized in Table 1. In this table, $\omega_{\text{rank}}$ refers to the exponent reported in Sedoglavic’s table (Sedoglavic, 2025), $\omega_{\text{new}}$ refers to the exponent obtained via Theorem 2, $L$ is the leading coefficient given by (2) and $A$ the number of additions as determined by the software of Mårtensson and Wagner  (Mårtensson and Wagner, 2025), and the structure indicator $1^{u}2^{v}3^{w}4^{x}$ means that we have found a decomposition consisting of $u$ copies of $\langle 1\rangle$, altogether $v$ tensors of rank 2 (i.e., $\langle 1,1,2\rangle$, $\langle 1,2,1\rangle$, or $\langle 2,1,1\rangle$), altogether $w$ tensors of rank 3 (a permutation of $\langle 1,1,3\rangle$), and $x$ tensors of rank 4 (either a permutation of $\langle 1,1,4\rangle$ or a permutation of $\langle 1,2,2\rangle$).
@@
More critically, we must take care that the structure imposed on the decomposition during steps 1 and 2 is preserved during lifting. This is not automatically ensured. For example, the decomposition of $\langle 6,6,6\rangle$ presented by Moosbauer and Poole contains two copies of $\langle 1,1,1\rangle$ which become a copy of $\langle 1,1,2\rangle$ when coefficients are taken modulo $2$.
@@ 6. Simulation
Since the bounds on the leading coefficient do not fully reflect the actual operation count of the algorithms, we simulate the recursive calls performed by our algorithm for different input sizes $N$ in order to get a more accurate idea of the runtime of the algorithm. Of course this approach does not take into account memory access costs and other practical considerations, but it does give a better estimate of the operation count than just the leading coefficient.
We compare Winograd’s variant of Strassen’s algorithm, which has a leading coefficient of $6$, to our algorithm using the decomposition of $\langle 6,6,6\rangle$ we found. The blue $\times$ and orange ++ in Figure 1 show the simulated operation counts for the algorithms, where for Moosbauer-Poole, we switch to Winograd-Strassen for $N<10^{4}$ and in both cases switch to the standard algorithm for $N<35$ as one would do in an actual implementation. The lines show the complexity estimate using the leading coefficient according to the formulas given above. The reason that our simulation shows lower operation counts than the reported bounds is that we switch to more efficient algorithms for small matrix sizes, while in the analysis we assumed that the recursion is performed exactly all the way down to $1\times 1$ matrices.

### v2

@@ Abstract.
journalyear: The results in this paper are flawed. Please do not use them. We are working on a revised version. conference: ; ; $c$s: Computing methodologies Algebraic algorithms
@@ 1. Introduction
We ran a search for such algorithms, first using a flip graph search to find algorithms using a minimal number of multiplications while retaining the special structure needed to apply our new recursive method. Then we optimize the number of additions using DeGroote actions. These steps produce an algorithm that is valid only modulo 22, thus we apply Hensel lifting to lift the algorithm to work over the integers, and finally we reduce the number of additions further by Mårtensson and Wagner’s common subexpression elimination method (Mårtensson and Wagner, 2025). We find improvements in the exponent for several small base cases including $3\times 3$ matrix multiplication, where we reduce the exponent from Laderman’s $2.854$ (Laderman, 1976) to $2.836$.
@@ 2. Background
Strassen’s algorithm (STRASSEN, 1969) uses a divide and conquer strategy to multiply matrices faster than the standard algorithm. First, it reduces the multiplication of two $n\times n$ matrices to multiplying two $2\times 2$ matrices whose entries are $\frac{n}{2}\times\frac{n}{2}$ matrices using block matrix multiplication. Then it computes this $2\times 2$ block matrix product using only 77 multiplications of the subblocks by linearly combining them in a clever way. Using this strategy recursively gives an algorithm to multiply $n\times n$ matrices using $\operatorname{O}(n^{\log_{2}(7)})$ operations in $R$. The exponent only depends on the number of multiplications, since the cost of the recursive call dominates the cost of the linear combinations.
@@
where $\mathbf{A}\in A,\mathbf{B}\in B$, $\mathbf{A}_{ij},\mathbf{B}_{ij}$ represent the $(i,j)$-th entry of $\mathbf{A}$ and $\mathbf{B}$ respectively and $\varGamma_{ij}\in C$ is the matrix with all zeros except for a 11 at the $(i,j)$-th entry. Note the $c_{ij}$ corresponds to $\mathbf{C}_{ji}$ in the product matrix; this is a standard convention to keep the cyclic nature of matrix multiplication.
@@
Using the decomposition Strassen found, we get $\langle 2,2,2\rangle\leq\langle 7\rangle$ and hence $\omega\leq\log_{2}(7)$. Further analysis allows to compute the leading coefficient of this algorithm as well, as we will see next. While the exponent is given by the rank of a decomposition, i.e. the number of multiplications, the leading coefficient depends on the number of additions and scalar multiplications needed in the restriction. We will denote this number by $A(\phi)$ if the restriction is given by $\phi$. The homomorphism for a restriction is in general not unique, but in our case every restriction is witnessed by an explicit homomorphism, so we just write AA when the homomorphism is clear from the context.
@@
We will show by induction on NN that
@@
In the case of Strassen’s original algorithm, which used a restriction $\langle 2,2,2\rangle\leq_{\phi}\langle 7\rangle$ with $A(\phi)=18$, we find that this theorem gives us a bound on the leading coefficient $L\leq 40$. Usually the leading coefficient bounds are computed in the idealized case where every recursive call is assumed to be done exactly. In this case, the leading coefficient of Strassen’s algorithm is 77, the general formula being $L\leq\frac{A}{r-n^{2}}+1$ (Karstadt and Schwartz, 2017).
@@
While we have to compute $\langle 1,1,2\rangle,\langle 1,2,2\rangle$ and their cyclic permutations by the standard algorithm, using 22 and 44 multiplications respectively, we can compute $\langle 2,2,2\rangle$ using Strassen’s algorithm with 77 multiplications. In the language of tensor restrictions we use $\langle 1,1,2\rangle\leq\langle 2\rangle$, $\langle 1,2,2\rangle\leq\langle 4\rangle$ and $\langle 2,2,2\rangle\leq\langle 7\rangle$, to conclude that $\langle 216,216,216\rangle\leq\langle 3581065\rangle$. Applying Theorem 3 to this restriction gives an algorithm with exponent $\omega_{0}=2.80751$, which is only slightly smaller than the original exponent $2.80754$. We can do even better by taking higher tensor powers.
@@ 4. Complexity Analysis
We will prove by induction on NN that
@@
We can count the arithmetic cost by considering performing this restriction recursively: we perform AA arithmetic operations on matrices of size $n^{k-1}\times n^{k-1}$ at the first step, so the arithmetic operation count at this step is bounded by $An^{2(k-1)}$. At the second step we have at most $r$ lots of $n^{k-2}\times n^{k-2}$ matrices that we need to construct, and then use linear combinations to get outputs; this means the arithmetic operations at this recursive step is bounded by $Arn^{2(k-2)}$. Continuing this, we see that at the $j$-th recursive step, we have at most $Ar^{j-1}n^{2(k-j)}$ arithmetic operations. Summing all of these, we get that the arithmetic cost is bounded by the geometric sum
@@ 5. Specific Decompositions
We have already used the Moosbauer and Poole’s tensor decomposition of $\langle 6,6,6\rangle$ in the form $137\odot\langle 1\rangle\oplus 8\odot\langle 2,1,1\rangle$ as illustrating example. According to Theorem 2, this decomposition leads to an exponent $2.80496$, slightly better than the exponent $2.80754$ announced by Moosbauer and Poole and than Strassen’s exponent $2.80735$, though not as good as the exponent $2.7925$ of the decompositions of Novikov et al. (Novikov et al., 2025b) and Dumas, Pernet, and Sedoglavic (Dumas et al., 2025). We have repeated the computation of Moosbauer and Pool in order to generate further decompositions of $\langle 6,6,6\rangle$ in the hope to find some that contain more copies of $\langle 2,1,1\rangle$. The best we found contains $18$ such copies and thus leads to an exponent $2.8016$. The leading coefficient for this decomposition turns out to be $7.22$, compared to 66 for Strassen’s algorithm.
@@
In a next step, for various choices of $n,m,p$, we have searched for decompositions of $\langle n,m,p\rangle$ that contain copies of $\langle 1,2,2\rangle$ and/or $\langle k,1,1\rangle$ for $k=2,3,4$, or permutations of these. This search procedure is described in more detail below. The results are summarized in Table 1. In this table, $\omega_{\text{rank}}$ refers to the exponent reported in Sedoglavic’s table (Sedoglavic, 2025), $\omega_{\text{new}}$ refers to the exponent obtained via Theorem 2, $L$ is the leading coefficient given by (2) and AA the number of additions as determined by the software of Mårtensson and Wagner  (Mårtensson and Wagner, 2025), and the structure indicator $1^{u}2^{v}3^{w}4^{x}$ means that we have found a decomposition consisting of $u$ copies of $\langle 1\rangle$, altogether $v$ tensors of rank 2 (i.e., $\langle 1,1,2\rangle$, $\langle 1,2,1\rangle$, or $\langle 2,1,1\rangle$), altogether $w$ tensors of rank 3 (a permutation of $\langle 1,1,3\rangle$), and $x$ tensors of rank 4 (either a permutation of $\langle 1,1,4\rangle$ or a permutation of $\langle 1,2,2\rangle$).
@@
More critically, we must take care that the structure imposed on the decomposition during steps 1 and 2 is preserved during lifting. This is not automatically ensured. For example, the decomposition of $\langle 6,6,6\rangle$ presented by Moosbauer and Poole contains two copies of $\langle 1,1,1\rangle$ which become a copy of $\langle 1,1,2\rangle$ when coefficients are taken modulo 22.
@@ 6. Simulation
Since the bounds on the leading coefficient do not fully reflect the actual operation count of the algorithms, we simulate the recursive calls performed by our algorithm for different input sizes NN in order to get a more accurate idea of the runtime of the algorithm. Of course this approach does not take into account memory access costs and other practical considerations, but it does give a better estimate of the operation count than just the leading coefficient.
We compare Winograd’s variant of Strassen’s algorithm, which has a leading coefficient of 66, to our algorithm using the decomposition of $\langle 6,6,6\rangle$ we found. The blue $\times$ and orange ++ in Figure 1 show the simulated operation counts for the algorithms, where for Moosbauer-Poole, we switch to Winograd-Strassen for $N<10^{4}$ and in both cases switch to the standard algorithm for $N<35$ as one would do in an actual implementation. The lines show the complexity estimate using the leading coefficient according to the formulas given above. The reason that our simulation shows lower operation counts than the reported bounds is that we switch to more efficient algorithms for small matrix sizes, while in the analysis we assumed that the recursion is performed exactly all the way down to $1\times 1$ matrices.